Describe how the x- and y-coordinates of the given quadratic functions differ from the x- and y-coordinates of corresponding points of y=x^2

y=(x+7)^2
1
Expert's answer
2020-12-10T07:44:11-0500


"y=(x+7)^2"

"=(x+7)(x+7)"

"=x(x+7)+7(x+7)"

"=x^2+7x+7x+49"

"=x^2+14x+49"

"a=1, b=14, and c=49"

"x=-b\/2a"

"=-14\/2"

"x=-7"

"y=(x)^2"

"y=(-7)^2"

"=49"

"Coordinates (-7, 49)"